The body, weighing 2 kg, moves rapidly the square root of 1.5 m / s at the point x = 1 m. The force acting on the body changes with location. a) Find the work done by the force on the body, when it moves from x = 1 m to x = 4 m. b) What is the velocity of the body at x = 4?


1
Expert's answer
2021-04-13T06:38:30-0400

Given:

"a=1.5\text{ m/s}^2,\\
m=2 \text{ kg},\\
v(x_1)=v_1=0."


Solution:

Find the force acting on the body:


"F=ma."

a) Find the work done by the force:


"W=Fd=F(x_2-x_1)=ma(x_2-x_1)=9\text{ J}."

b) The velocity is


"v_4=\sqrt{2ad}=\sqrt{2a(x_2-x_1)}=3\text{ m/s}."